• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name M Jones

    Log entry time 08:05:36 on October26,2002

    Entry number 89040

    keyword=owl shift summary

    MCC is trying to give us beam.
    We tried to do a harp scan, but the software failed. See halog 89020.
    Starting around 00:00 MCC is trying to give max.
    Right now 00:10 we are getting 10uA . Start run just to checkout
    Have about 50uA.
    Run 2750
    Looked at count rate for 2750. It is the same as 2748 within errors.
    MCC can't get 0,0 at 4B and energy fast feedback lock.
    Opt to stay with energy lock and beam at 0.6 ,-0.4 at 4B.
    Start running at 100uA about 1:45.
    Having been running steady at kin q3d_e40 .